# Integrated Communication and Bayesian Estimation of Fixed Channel States

### v1

@@ III Main Result
$\operatorname{\sf E}^{2}[J_{{\sf DP},x^{n}}^{-1}(S)]=\frac{1}{n^{2}}\left(\operatorname{\sf E}_{S}^{2}[\operatorname{\sf E}_{X}[J_{{\sf D},X}^{-1}(s)]]+o(1)\right),$
@@
$=\frac{1}{n}\left(\operatorname{\sf E}_{S}[\operatorname{\sf E}_{X}[J_{{\sf D},X}^{-1}(s)]]+o(1)\right)+\frac{\text{const}}{n^{2}}+\frac{\text{const}}{n^{2}},$
@@
$n\operatorname{\sf E}[(S-{\hat{s}}(Y^{n}|x^{n}))^{2}]\to\operatorname{\sf E}_{S}[\operatorname{\sf E}_{X}[J_{{\sf D},X}^{-1}(s)]].$
@@
$\alpha\leq\operatorname{\sf E}_{S}[\operatorname{\sf E}_{X}[J_{{\sf D},X}^{-1}(s)]].$
@@
${\sf MSE}({\hat{s}}|x^{n})\geq\frac{\operatorname{\sf E}_{S}[\operatorname{\sf E}_{X}[J_{{\sf D},X}^{-1}(s)]]+o(1)}{n}.$

### v2

@@ III Main Result
$\operatorname{\sf E}^{2}[J_{{\sf DP},x^{n}}^{-1}(S)]=\frac{1}{n^{2}}\left(\operatorname{\sf E}_{S}^{2}[\operatorname{\sf E}_{X}^{-1}[J_{{\sf D},X}(s)]]+o(1)\right),$
@@
$=\frac{1}{n}\left(\operatorname{\sf E}_{S}[\operatorname{\sf E}_{X}^{-1}[J_{{\sf D},X}(s)]]+o(1)\right)+\frac{\text{const}}{n^{2}}+\frac{\text{const}}{n^{2}},$
@@
$n\operatorname{\sf E}[(S-{\hat{s}}(Y^{n}|x^{n}))^{2}]\to\operatorname{\sf E}_{S}[\operatorname{\sf E}_{X}^{-1}[J_{{\sf D},X}(s)]].$
@@
$\alpha\leq\operatorname{\sf E}_{S}[\operatorname{\sf E}_{X}^{-1}[J_{{\sf D},X}(s)]].$
@@
${\sf MSE}({\hat{s}}|x^{n})\geq\frac{\operatorname{\sf E}_{S}[\operatorname{\sf E}_{X}^{-1}[J_{{\sf D},X}(s)]]+o(1)}{n}.$